If two whole numbers have a sum of 18 and a product of 45, what are the numbers? (32) Place the numbers 1 through 9 in the cells so that the sum in each direction is 15.

Answer by ankor@dixie-net.com(22740)

If two whole numbers have a sum of 18 and a product of 45, what are the numbers?
two equations
a + b = 18
b = (18-a)
and
a * b = 45
Replace b with (18-a)
a*(18-a) = 45
18a - a^2 = 45
0 = a^2 - 18a + 45
Factors to
(a-15)(a-3) = 0
Two solutions
a = 15, then b=3
and
a = 3, then b=15

(32) Place the numbers 1 through 9 in the cells so that the sum in each direction is 15.
How about
1 8 6
9 4 2
5 3 7